# Query-Level Uncertainty in Large Language Models

### v3

@@ 4.3 Internal Confidence is Much Faster than Answer-Level Approaches
We compare our query-level Internal Confidence with several popular answer-level uncertainty methods on GSM8K using Qwen-14B, including Perplexity (Fomicheva et al., 2020), Semantic Entropy (Kuhn et al., 2023), $\textrm{P}(\text{True})$ (Kadavath et al., 2022), Lexical Similarity (Fomicheva et al., 2020), and SAR (Duan et al., 2024).
Method $\uparrow$ AUC $\downarrow$ Time (s) $\uparrow$ Speedup
@@
Lexical Similarity 62.4 22.3 $74\times$
@@
Table 2: Comparison of query-level Internal Confidence with answer-level uncertainty methods (Qwen-14B on GSM8K).
Table 2 compares the effectiveness and runtime across different approaches. While answer-level approaches such as Perplexity, $\textrm{P}(\text{True})$, and SAR require significantly higher computation time (ranging from nearly 10 seconds up to more than 180 seconds per sample), our Internal Confidence method achieves the best AUC (66.8) with an average running time of only 0.3 seconds. This corresponds to speedups of over 30× to 600× compared to existing baselines. These results demonstrate that Internal Confidence combines state-of-the-art accuracy with an extremely fast inference speed, which can be a practical choice for large-scale or latency-sensitive reasoning tasks.
Notably, the running time for Internal Confidence remains constant, independent of the length of answers. Figure 5 shows that the runtime of the best answer-level approach, SAR, grows with the answer length, reaching nearly 500s for answers over 600 characters. In contrast, Internal Confidence achieves large acceleration ratios (736×–1672×), with speedups increasing as answers become longer, which demonstrates its scalability and efficiency. See results of other datasets in Table A1.
@@ 4.4 Internal Confidence Makes LLM Reasoning More Efficient
We conduct experiments for two scenarios: (1) Efficient RAG. Basically, the Internal Confidence can serve as a signal of the knowledge gaps of a model. If the score is greater than a threshold, the model is confident to address the query. Otherwise, it requires the call of RAG. We use the TriviaQA dataset for evaluation. This dataset provides web search results for a query, which can be used as retrieved contexts for RAG. (2) Model Cascading. This task aims to achieve cost-performance trade-offs by coordinating small and large models (Dohan et al., 2022; Gupta et al., 2024). The smaller models are responsible for easy assignments. If they are aware that the mission is hard to complete, they invoke a larger model. We use a two-model cascade setting with Phi-3.8B and Llama-8B on the TriviaQA dataset. If the Internal Confidence of the smaller model is high, we do not invoke the larger model. Otherwise, the hard query is deferred to the larger model.
@@ 4.5 Locality Affects Uncertainty Performance
Our method incorporates attenuated encodings to aggregate probabilities centering around a decision point. The locality of the encoding may affect the accuracy of estimated uncertainties. To study the influence of the locality, we vary the $w$ in Equation 4 to obtain encodings with different localities and observe how they affect the estimations. Figure 5 reports the average AUC across three datasets and models. The results indicate that the effect of locality depends on both the task type and the model architecture. Although the optimal locality may vary with model and dataset (see details in Section C.3), we find that a default setting of $w=1.0$ (corresponding to Locality $\approx 0.7$) yields consistently competitive performance that generalize well.
@@ A.1 Aleatoric and Epistemic Uncertainty
Uncertainty in machine learning is commonly categorized into two main types: aleatoric and epistemic uncertainty (Hora, 1996; Der Kiureghian & Ditlevsen, 2009; Hüllermeier & Waegeman, 2021). These distinctions are often overlooked in the context of LLM uncertainty estimation. Aleatoric uncertainty arises from inherent randomness in the data, such as ambiguous inputs or conflicting annotations. This type of uncertainty is irreducible, as it reflects intrinsic noise in the input data. In contrast, epistemic uncertainty stems from a lack of knowledge, often due to insufficient training data and limited model capacity. Unlike aleatoric uncertainty, epistemic uncertainty is reducible with additional data or advanced modeling. In this work, we focus specifically on epistemic uncertainty, with the goal of evaluating whether an LLM possesses sufficient knowledge to answer a given query. For evaluation, we adopt factual QA and mathematical reasoning benchmarks, which are designed to have clear-cut answers. We assume these datasets are well-curated to minimize aleatoric uncertainty, such as ambiguous questions and inconsistent labels. However, we acknowledge that residual ambiguity may persist, given the inherent nature of linguistic ambiguity (Gillon, 1990) and the difficulty of fully disentangling aleatoric from epistemic uncertainty (Mucsányi et al., 2024). We treat such aleatoric effects as negligible for the purposes of focusing on epistemic uncertainty.
@@ A.2 Uncertainty and Confidence
Appendix B Baseline Details
We adapt existing answer-level methods to quantify the pre-generation uncertainty, e.g., logit-based uncertainty. Given a query (including the prompt) $\mathbf{x}=(x_{1},\ldots,x_{N})$, we can obtain a probability for each token $P(x_{n}\mid x_{<n})$ by performing a forward pass. (1) The baseline $\text{Max}(-\log p)$ measures the query’s uncertainty by assessing the least likely token in the query (Manakul et al., 2023). (2) Predictive Entropy is defined as the entropy over the entire query token sequence (Malinin & Gales, 2021):
@@ Appendix B Baseline Details
(3) Min-K Entropy combines the ideas of $\text{Max}(-\log p)$ and predictive entropy, by selecting the top-$K$ tokens from the query with the minimum token probability (Shi et al., 2024). (4) Attentional Entropy is a modified version of the predictive entropy that considers a weighted sum:
@@
Method $\uparrow$ AUC $\downarrow$ Time (s) $\uparrow$ Speedup
@@
(a) TriviaQA
(b) SciQ
(c) GSM8K
Figure A1: Distinguishing between known and unknown queries using Internal Confidence for Phi-3.8B.
@@
Appendix C Additional Experiments
C.1 Calibration Performance
Figure A1 compares the distributions of Internal Confidence scores for known (green) and unknown (blue) queries across three datasets. The results reveal that Internal Confidence tends to assign higher values to known queries and lower values to unknown queries, which is suitable for distinguishing the two groups. Specifically, on TriviaQA, the separation is mild with noticeable overlap. On SciQ, the known queries concentrate near 1.0, while unknown queries spread toward lower scores, and on GSM8K, the distinction is the clearest, with known queries clustered in the high-confidence region (0.8–0.9) and unknown queries shifted leftward.
@@ C.1 Calibration Performance
C.2 Internal Confidence does not rely on in-context learning
Figure A2 shows the effect of the number of in-context learning example pairs ($k$-shot) on model performance across three datasets and models. Here, we randomly select $k$ pairs of positive and negative samples. We plot the AUC as a function of $k$-shot values from 1 to 5. Overall, Llama-8B and Qwen-14B maintain relatively stable performance with slight improvements as $k$ increases, while Phi-3.8B exhibits more fluctuation, especially on TriviaQA. These results suggest that the benefit of additional in-context examples varies across both models and datasets. Therefore, our Internal Confidence can obtain strong performance even without in-context learning from examples, which can reduce the computational cost.
C.3 Impact of Locality
Figure A3 presents the impact of locality on AUC performance across three datasets (TriviaQA, SciQ, GSM8K) and three models (Phi-3.8B, Llama-8B, Qwen-14B). For Phi-3.8B, AUC improves gradually with increasing locality across all datasets, with TriviaQA exhibiting consistently higher discriminability than SciQ and GSM8K. For Llama-8B, the performance remains fairly stable across different locality values, showing only minor fluctuations, particularly for SciQ and GSM8K. For Qwen-14B, the AUC increases with the locality for all datasets up to a certain point, after which it either plateaus or slightly declines; this trend is most evident for GSM8K.
@@ C.3 Impact of Locality
Appendix D Use of Large Language Models

### v4

@@ 4.3 Internal Confidence is Much Faster than Answer-Level Approaches
We compare our query-level Internal Confidence with several popular answer-level uncertainty methods on GSM8K using Qwen-14B, including Perplexity (Fomicheva et al., 2020), Semantic Entropy (Kuhn et al., 2023), $\textrm{P}(\text{True})$ (Kadavath et al., 2022), Lexical Similarity (Fomicheva et al., 2020), SAR (Duan et al., 2024), Maximum Sequence Probability (MSP), CCP (Fadeeva et al., 2024), and EV Laplacian (Lin et al., 2023).
Method $\uparrow$ AUROC $\downarrow$ Time (s) $\uparrow$ Speedup
@@
Lexical Similarity 62.4 170.3 $567\times$
@@
Table 2: Comparison with answer-level uncertainty methods (Qwen-14B on GSM8K).
Table 2 compares the effectiveness and runtime across different approaches. While answer-level approaches such as Perplexity, $\textrm{P}(\text{True})$, and SAR require significantly higher computation time (ranging from nearly 10 seconds up to more than 180 seconds per sample), our Internal Confidence method achieves the best AUROC (66.8) with an average running time of only 0.3 seconds. This corresponds to speedups of over 30× to 600× compared to existing baselines. These results demonstrate that Internal Confidence achieves competitive performances compared to answer-level uncertainty approaches while being extremely faster, which can be a practical choice for tasks requiring longer and more complex answers.
Notably, the running time for Internal Confidence remains constant, independent of the length of answers. Figure 5 shows that the runtime of the best answer-level approach, SAR, grows with the answer length, reaching nearly 500s for answers over 600 characters. In contrast, Internal Confidence achieves large acceleration ratios, with speedups increasing as answers become longer, which demonstrates its scalability and efficiency. See results of other datasets in Table A1.
@@ 4.4 Internal Confidence Makes LLM Reasoning More Efficient
We conduct experiments for two scenarios: (1) Efficient (or adaptive) RAG. This assumes that the Internal Confidence can serve as a signal of the knowledge gaps of a model. If the score is greater than a threshold, the model is sufficiently confident to address the query directly using its parametric knowledge. Otherwise, it requires invoking RAG. Existing studies have explore adaptive RAG through learned classifiers (Jeong et al., 2024; Marina et al., 2025) and answer-level uncertainty approaches (Jiang et al., 2023; Su et al., 2024; Yao et al., 2025; Moskvoretskii et al., 2025), which actively decide whether and when to retrieve documents. However, these approaches require training samples or generating answers to measure the uncertainty. In contrast, our Internal Confidence method is training-free and significantly faster than answer-level approaches (as shown in Table 2), which can serve as a potentially efficient way to guide adaptive RAG. We use the TriviaQA dataset for evaluation. This dataset provides web search results for a query, which can be used as retrieved contexts for RAG. (2) Model Cascading. This task aims to achieve cost-performance trade-offs by coordinating small and large models (Dohan et al., 2022; Gupta et al., 2024). The smaller models are responsible for easy assignments. Whenever they determine that the task exceeds their capabilities, it can be delegated to a larger model. We use a two-model cascade setting with Phi-3.8B and Llama-8B on the TriviaQA dataset. If the Internal Confidence of the smaller model is high, we do not invoke the larger model. Otherwise, the hard query is deferred to the larger model.
@@
Figure 5: Impact of locality on validation set performance. We report the average AUROC across the three considered datasets. See details in Section D.3.
@@ 4.5 Locality Affects Uncertainty Performance
Our method incorporates attenuated encodings to aggregate probabilities centering around a decision point. The locality of the encoding may affect the accuracy of estimated uncertainties. To study the influence of the locality, we vary the $\alpha$ in Equation 4 to obtain encodings with different localities and observe how they affect the estimations. Figure 5 reports the average AUROC across three datasets and models. The results indicate that the effect of locality depends on both the task type and the model architecture. Although the optimal locality may vary with model and dataset (see details in Section D.3), we find that a default setting of $\alpha=1.0$ (corresponding to Locality $\approx 0.7$) yields consistently competitive performance that generalize well.
@@ 5 Conclusion
Limitations
There are several main limitations of this work. (1) Our proposed query-level uncertainty measure relies on access to a model’s internal states, which is not feasible for fully black-box APIs. (2) We adopt straightforward fixed hyperparameters across all experiments for efficiency and generalizability, but this choice does not yield optimal performance in all settings. As discussed in Section D.5, our additional experiments show that the optimal decision center location varies across models and tasks. (3) Although internal confidence can serve as a strong baseline for detecting knowledge boundary, its performance still lags behind answer-level approaches. We hope this work inspires future research on more refined and robust ways to detect the knowledge boundary of foundation models.
@@ A.1 Aleatoric and Epistemic Uncertainty
Uncertainty in machine learning is commonly categorized into two main types: aleatoric and epistemic uncertainty (Hora, 1996; Der Kiureghian and Ditlevsen, 2009; Hüllermeier and Waegeman, 2021). These distinctions are often overlooked in the context of LLM uncertainty estimation. Aleatoric uncertainty arises from inherent randomness in the data, such as ambiguous inputs or conflicting annotations. This type of uncertainty is irreducible, as it reflects intrinsic noise in the input data. In contrast, epistemic uncertainty stems from a lack of knowledge, often due to insufficient training data and limited model capacity. Unlike aleatoric uncertainty, epistemic uncertainty is reducible with additional data or advanced modeling. In this work, we focus specifically on epistemic uncertainty, with the goal of evaluating whether an LLM possesses sufficient knowledge to answer a given query. For evaluation, we adopt factual QA and mathematical reasoning benchmarks, which are designed to have clear-cut answers. We assume these datasets are well-curated to minimize aleatoric uncertainty, such as ambiguous questions and inconsistent labels. However, we acknowledge that residual ambiguity may persist, given the inherent nature of linguistic ambiguity (Gillon, 1990) and the difficulty of fully disentangling aleatoric from epistemic uncertainty (Mucsányi et al., 2024). We treat such aleatoric effects as negligible for the purposes of focusing on epistemic uncertainty.
@@ A.2 Uncertainty and Confidence
(a) TriviaQA
(b) SciQ
(c) GSM8K
Figure A1: Distinguishing between answerable and non-answerable queries using Internal Confidence for Phi-3.8B.
Appendix B Prompt
We use the following prompt template for all experiments. The query $x$ consists of both prompt and question tokens.
You are a helpful assistant that assesses whether you can provide an accurate response to a question. Respond only with ’Yes’ or ’No’ to indicate whether you are capable of answering the following question. {Examples}{Input Question}.
Appendix C Baseline Details
We adapt existing answer-level methods to quantify the pre-generation uncertainty, e.g., logit-based uncertainty. Given a query (including the prompt) $\mathbf{x}=(x_{1},\ldots,x_{N})$, we can obtain a probability for each token $P(x_{n}\mid x_{<n})$ by performing a forward pass. (1) The baseline $\text{Max}(-\log p)$ measures the query’s uncertainty by assessing the least likely token in the query (Manakul et al., 2023). (2) Predictive Entropy is defined as the entropy over the entire query token sequence (Malinin and Gales, 2021):
@@ Appendix C Baseline Details
(3) Min-K Entropy combines the ideas of $\text{Max}(-\log p)$ and predictive entropy, by selecting the top-$K$ tokens from the query with the minimum token probability (Shi et al., 2024). (4) Attentional Entropy is a modified version of the predictive entropy that considers a weighted sum
@@
Method $\uparrow$ AUROC $\downarrow$ Time (s) $\uparrow$ Speedup
@@
Appendix D Additional Experiments
D.1 Calibration Performance
Figure A1 illustrates the distributions of Internal Confidence for answerable versus non-answerable queries across three datasets—TriviaQA, SciQ, and GSM8K—using Phi-3.8B. In all cases, answerable queries (green) exhibit noticeably higher Internal Confidence, with distributions concentrated toward the upper end of the confidence range. In contrast, non-answerable queries (blue) show substantially lower Internal Confidence, typically forming broader or left-shifted distributions. This clear separation demonstrates that Internal Confidence effectively distinguishes between seen and unseen inputs, supporting its usefulness as an internal signal for assessing familiarity and reliability within the model.
@@ D.1 Calibration Performance
D.2 Internal Confidence does not rely on in-context learning
Figure A2 shows the effect of the number of in-context learning example pairs ($k$-shot) on model performance across three datasets and models. Here, we randomly select $k$ pairs of positive and negative samples. We plot the AUROC as a function of $k$-shot values from 1 to 5. Overall, Llama-8B and Qwen-14B maintain relatively stable performance with slight improvements as $k$ increases, while Phi-3.8B exhibits more fluctuation, especially on TriviaQA. These results suggest that the benefit of additional in-context examples varies across both models and datasets. Therefore, our Internal Confidence can obtain strong performance even without in-context learning from examples, which can reduce the computational cost.
D.3 Impact of Locality
Figure A3 presents the impact of locality on AUROC performance across three datasets (TriviaQA, SciQ, GSM8K) and three models (Phi-3.8B, Llama-8B, Qwen-14B). For Phi-3.8B, AUROC improves gradually with increasing locality across all datasets, with TriviaQA exhibiting consistently higher discriminability than SciQ and GSM8K. For Llama-8B, the performance remains fairly stable across different locality values, showing only minor fluctuations, particularly for SciQ and GSM8K. For Qwen-14B, the AUROC increases with the locality for all datasets up to a certain point, after which it either plateaus or slightly declines; this trend is most evident for GSM8K.
@@ D.3 Impact of Locality
D.4 Internal Confidence Can Be Generalized to More Challenging Datasets
To validate whether our proposed internal confidence can be generalized to more challenging tasks, we conduct experiments on three additional datasets: (1) SimpleQA (Wei et al., 2024). This is a benchmark that evaluates the ability of language models to answer short, fact-relevant questions, which is less likely to be contaminated by the pre-training stage. (2) MuSiQue (Trivedi et al., 2022). This is a dataset that requires proper multihop reasoning, which is more difficult and harder to cheat via disconnected reasoning. (3) TruthfulQA (Lin et al., 2022). This is a benchmark to measure whether a language model is truthful in generating answers to questions. The authors crafted questions that some humans would answer falsely due to a false belief or misconception. To perform well, a model has to avoid generating false answers learned from imitating human texts. For each dataset, we use the validation or test partition for comparison, which contains a reasonable number of samples (2-4K). For the first two datasets, we apply the default configuration of internal confidence. Regarding the TruthfulQA dataset, we observe that the task exhibits a distinct decision center, which tends to appear in the middle layers rather than the upper layers across all three model architectures. For example, on a 100-sample held-out validation set, the decision centers appear at layers 9, 7, and 23 for Phi-3.8B, Llama-8B, and Qwen-14B, respectively. To consider this, we learn the decision center specifically for TruthfulQA using a 100-sample validation set. The overall results are shown in Table A2. We can observe that our internal confidence can outperform other query-level uncertainty consistently across datasets and architectures.
D.5 The Optimal Decision Center Varies Across Models and Tasks
We use the top right position as a default decision center, which offers a training-free and pragmatic solution, but it is the optimal center. We conduct experiments to study the learned decision center across different models and tasks. Figure A4, Figure A5, and Figure A6 show the locations of decision centers. We can observe that the center tends to appear at the top right place for TriviaQA and SciQ while the math reasoning task of GSM8K has a distinct behavior. The center is located in the lower layers. Although the current default center (top right) is sub-optimal, it offers a training-free, strong baseline, which can be generalized across different applications.
(a) Phi-3.8B
(b) Llama-8B
(c) Qwen-14B
Figure A4: Learned decision centers of Phi-3.8B.
(a) Phi-3.8B
(b) Llama-8B
(c) Qwen-14B
Figure A5: Learned decision centers of Llama-8B.
(a) Phi-3.8B
(b) Llama-8B
(c) Qwen-14B
Figure A6: Learned decision centers of Qwen-14B.
Appendix E Use of Large Language Models
@@
SimpleQA MuSiQue TruthfulQA Avg
Method $\uparrow$ AUROC $\uparrow$ PRR $\downarrow$ ECE $\uparrow$ AUROC $\uparrow$ PRR $\downarrow$ ECE $\uparrow$ AUROC $\uparrow$ PRR $\downarrow$ ECE $\uparrow$ AUROC $\uparrow$ PRR $\downarrow$ ECE
Phi-3.8B
$\text{Max}(-\log p)$ 50.5 5.5 —- 53.2 4.7 —- 50.3 -1.3 —- 51.3 3.0 —-
Predictive Entropy 54.0 5.9 —- 65.6 29.6 —- 55.7 15.6 —- 58.4 17.0 —-
Min-K Entropy 53.8 13.2 —- 59.9 21.9 —- 55.1 13.9 —- 56.3 16.3 —-
Attentional Entropy 50.1 5.1 —- 54.7 11.6 —- 52.4 8.3 —- 52.4 8.3 —-
Perplexity 52.4 5.8 —- 55.2 7.1 —- 54.1 3.8 —- 53.9 5.6 —-
$\textrm{P}(\text{Yes})$ (top right) 61.3 17.8 18.8 65.4 29.8 9.5 39.6 -16.2 27.1 55.4 10.5 18.5
$\textrm{P}(\text{Yes})$ (naive avg) 59.8 17.8 69.9 65.5 29.5 63.2 49.3 -2.0 25.5 58.2 15.1 52.9
Internal Confidence 61.2 26.1 18.2 65.5 30.2 9.3 56.4 13.2 40.7 61.0 23.2 22.7
Llama-8B
$\text{Max}(-\log p)$ 50.1 -2.9 —- 53.2 6.3 —- 52.4 4.1 —- 51.9 2.5 —-
Predictive Entropy 49.1 -0.9 —- 56.4 13.1 —- 60.0 13.7 —- 55.2 8.6 —-
Min-K Entropy 49.8 0.1 —- 56.0 15.2 —- 57.8 21.0 —- 54.5 12.1 —-
Attentional Entropy 48.6 -4.2 —- 57.4 17.7 —- 53.3 9.6 —- 54.1 7.7 —-
Perplexity 50.1 -3.8 —- 54.2 5.9 —- 54.3 4.8 —- 52.9 2.3 —-
$\textrm{P}(\text{Yes})$ (top right) 53.6 5.2 78.9 64.1 27.3 74.5 43.3 -11.9 55.7 53.7 6.9 69.7
$\textrm{P}(\text{Yes})$ (naive avg) 54.9 5.8 28.5 63.2 31.9 18.5 47.0 1.9 3.4 55.0 13.2 16.8
Internal Confidence 55.6 11.6 67.4 64.3 29.8 74.4 63.2 26.8 15.4 61.0 22.7 52.4
Qwen-14B
$\text{Max}(-\log p)$ 50.8 -1.2 —- 52.5 6.8 —- 51.0 4.7 —- 51.4 3.4 —-
Predictive Entropy 50.5 1.3 —- 53.8 9.4 —- 59.9 21.1 —- 54.7 10.6 —-
Min-K Entropy 51.8 8.1 —- 54.1 2.3 —- 58.1 22.8 —- 54.7 11.1 —-
Attentional Entropy 48.7 -2.9 —- 54.9 13.2 —- 50.7 3.2 —- 51.4 4.5 —-
Perplexity 50.1 -2.4 —- 53.7 10.7 —- 52.6 6.0 —- 52.1 4.8 —-
$\textrm{P}(\text{Yes})$ (top right) 55.6 11.4 35.5 57.7 14.4 22.4 42.6 -19.0 53.4 52.0 2.3 37.1
$\textrm{P}(\text{Yes})$ (naive avg) 57.6 16.6 4.0 58.4 16.9 6.5 49.7 -3.5 8.6 55.2 10.0 6.4
Internal Confidence 56.0 13.2 10.3 57.6 14.6 5.4 58.0 16.4 0.5 57.2 14.7 5.4
Table A2: Additional results of different query-level uncertainty estimation methods.